# Recommendation Unlearning via Influence Function

### v2

@@ Abstract.
In this work, we propose a new Influence Function-based Recommendation Unlearning (IFRU) framework, which efficiently updates the model without retraining by estimating the influence of the unusable data on the model via the influence function. In the light that recent recommender models use historical data for both the constructions of the optimization loss and the computational graph (e.g., neighborhood aggregation), IFRU jointly estimates the direct influence of unusable data on optimization loss and the spillover influence on the computational graph to pursue complete unlearning. Furthermore, we propose an importance-based pruning algorithm to reduce the cost of the influence function. IFRU is harmless and applicable to mainstream differentiable models. Extensive experiments demonstrate that IFRU achieves more than 250 times acceleration compared to retraining-based methods with recommendation performance comparable to full retraining. Codes are avaiable at https://github.com/baiyimeng/IFRU.
Recommender System, Recommendation Unlearning, Privacy, Influence Function
copyright: acmlicensedjournal: TORSjournalyear: 2023journalvolume: xxxjournalnumber: xxxarticle: 1publicationmonth: 6price: 15.00doi: 1xxxxx/xxxxxxccs: Information systems Recommender systemsccs: Information systems Collaborative filteringccs: Security and privacy Privacy protections
@@
We propose a new Influence Function-based Recommendation Unlearning (IFRU) framework to achieve more complete unlearning in recommendation. The IFRU extends the influence function to quantify the influence of unusable data in the computational graph aspect, termed spillover influence. Specifically, IFRU expresses the computational graph change after erasing the unusable data in the optimization loss form, making spillover influence measurable like the direct influence. Taking one step further, we reduce the cost of influence function that computes the Hessian matrix with the size of model parameters by an importance-based pruning method to ignore less affected model parameters. IFRU avoids alterations to the training process and model architecture, emphasizing its capacity to maintain harmlessness while facilitating its application into mainstream models. We instantiate IFRU on two representative recommender models: MF (Koren et al., 2009) and LightGCN (He et al., 2020), and conduct experiments on two real-world datasets to demonstrate our method.
@@ 2.1. Problem Formulation
where $\hat{\theta}$ denotes the model parameters learned from $\mathcal{D}$, $L(\theta;\mathcal{D})$ denotes the total loss on $\mathcal{D}$, and $\ell(\cdot)$ is a recommendation loss function. In this study, our primary focus lies on the pointwise loss, particularly in investigating the Binary Cross-Entropy (BCE) loss. Nonetheless, our method can readily extend to other types of losses, such as pair-wise loss BPR (Rendle et al., 2012).
@@ 3. Methodology
In this section, we first introduce how IFRU achieves complete-efficient-harmless recomendation unlearning via the influence function (Section 3.1), and further acceleration via pruning (Section 3.2). We then present the instantiation of IFRU (Section 3.3) and further discussions of the proposal (Section 3.4).
@@ 3.1.1. Derivation of $\Delta$
Here $\hat{\theta}_{\epsilon}$ represents the optimal model parameters under this scenario.
@@ 3.2. Unlearning with Pruning
Input: Pruning ratios $\{a_{0},...,a_{K}\}$, hyper-parameters $K$, training data $\mathcal{D}$, and the unusable data $\mathcal{D}_{r}$.
1 For each $k\in\{0,1,\dots,K\}$, initialize $s_{v}^{k}$ as zero for each $v$ (a user or an item), and initialize $\mathcal{S}^{k}=\varnothing$ ;
@@ 3.3. Instantiation
Instantiation on LighGCN. For LightGCN, the computational graph for some remaining data will possibly change if removing $\mathcal{D}_{r}$, e.g., the aggregation operations could be different under $\mathcal{D}$ and $\mathcal{D}_{-r}$ as shown in Figure 5. That means, for some training data $(u,i,y_{ui})\in\mathcal{D}_{-r}$, we have $f_{\theta}(u,i;\mathcal{D})\neq f_{\theta}(u,i;\mathcal{D}_{-r})$. Thus $\mathcal{D}_{c}$ is not $\varnothing$. For simplicity, we set the number of graph convolution layers as $1$. In this case, it is easily verified that only the interaction (in $\mathcal{D}_{-r}$) of the users/items having interactions in $\mathcal{D}_{r}$ belongs to $\mathcal{D}_{c}$. That means:
@@ 3.4. Discussion
    As we consider both the spillover and direct influence, the estimated $\mathcal{I}(\hat{\theta};\mathcal{D}_{r})$ can faithfully reflect the true $\frac{d\hat{\theta}_{\epsilon}}{d\epsilon}\big{|}_{\epsilon=-\frac{1}{\mathcal{D}}}$. Then, if $|\mathcal{D}|$ is large enough, $o(\frac{1}{|\mathcal{D}|})$ is ignorable and $\hat{\theta}^{\prime}$ is a good approximation for the $\hat{\theta}^{*}$. Then, our unlearned model $f_{\hat{\theta}^{\prime}}(u,i;\mathcal{D}_{-r})$ can approach $f_{\hat{\theta}^{*}}(u,i;\mathcal{D}_{-r})$ well777 It is important to note that there are two underlying assumptions: 1) we assume that the errors caused by randomness and convergence in the learning models and calculating the influence can be ignored, and 2) it is assumed that the objective is strongly convex. Recent research by (Bae et al., 2022) suggests that the errors caused by randomness and convergence are very small. Furthermore, under non-convex cases, we can add the identity matrix with a damping term (Koh and Liang, 2017) to refine the defined influence, making it still applicable with smaller approximation errors (Bae et al., 2022; Warnecke et al., 2023). . Therefore, IFRU could achieve approximatively complete unlearning (defined in Section 2.1) if the size of training data is large enough. Regarding the pruning, empirical results (cf. Section 4.3.3) show that it will only slightly sacrifice the unlearning completeness since we only prune the model parameters less affected by erasing $\mathcal{D}_{r}$.
@@ 4.1.4. Hyper-parameters Setting
For fair comparisons, we use BCE loss (He et al., 2017) to learn model parameters for all methods [^10] Our codes will be made publicly available at the time of publication to ensure complete reproducibility of the experiments.. Meanwhile, we optimize all models using the Adam optimizer with a default batch size of 2048 and set the maximal number of epochs to 5000. The best hyper-parameters are determined based on the result of the validation set. The embeddings of the backbone recommender are initialized with the Gaussian distribution, the mean of which is fixed to zero and the standard deviation of which is tuned in the range of {1e-2, 1e-3, 1e-4}. The embedding size is tuned in the range of {32, 48, 64}. The early stopping strategy is applied to prevent the over-fitting problem, that is, terminating model training if the $AUC_{0}$ metric on the validation set does not increase in 50 successive epochs. The ratio of unusable data $\mathcal{D}_{r}$ to training data $\mathcal{D}$, i.e., $|\mathcal{D}_{r}|/|\mathcal{D}|$ is set to $2\%$ for MF and $1\%$ for LightGCN by default [^11] We would include a study that does not use HVP, where unlearning for LightGCN would be costly. Thus We take a smaller ratio for LightGCN to ensure the study could be easily conducted. For all the compared methods, we tune the learning rate in the range of {1e-2, 1e-3, 1e-4}. For RecEarser, we adopt the Interaction-based Balanced Partition algorithm for data splitting and set the number of shards to 10, as suggested by the authors in their paper. The number of shards in SISA is also set to 10. For IFRU, we fix $K$ in Algorithm 1 to 1, as $1$-hop neighbors cover almost all users and items under the default ratio of unusable data. Then, we set the pruning ratios $(a_{0},a_{1})$ of Algorithm 1 to (1.0, 1.0) for MF and (1.0, 0.6) for LightGCN by default. The learning rate of Adam optimizer to calculate the defined influence is tuned in the range of {1e3, 2e3, 5e3, 1e4, 2e4, 5e4}. Notably, we have also tried smaller values for this learning rate and larger values for other learning rates. However, both of them lead to much worse results.
@@ 4.2.1. Accuracy Comparison
Harmless unlearning. Both incomplete unlearning and not achieving harmless unlearning can result in performance gaps between the Retrain method and unlearning methods, which may raise doubts about the superiority of IFRU in achieving harmless unlearning. To provide direct evidence for the superiority, we conduct an additional experiment. Specifically, after the backbone model was trained with a dataset, we forcibly set the unusable data $\mathcal{D}_{r}$ with an empty set and performed unlearning on it. In this case, if there is still a performance gap to Retrain, it must be due to not achieving harmless unlearning, as removing an empty set would not bring parameter changes for any method, including IFRU (seeing Equation (13)). The results of the experiment are summarized [^13] Notably, the training datasets for MF and LightGCN are $\mathcal{D}$ and $\mathcal{D}_{-r}$, respectively. Utilizing $\mathcal{D}_{-r}$ leads to performance discrepancies between Figure 4 and Table 2 for LightGCN. in Figure 4. The figure shows that IFRU/SCIF achieves the same performance as Original and Retrain, confirming that unlearning with influence function can help achieve harmless unlearning. In contrast, RecRraser and SISA show performance gaps to Original and Retrain, indicating that harmless unlearning is not achieved.
@@ 5.1. Machine Unlearning
In addition to researching how to achieve unlearning, there are also some works that specifically focus on the benchmarks (Schelter, [n. d.]), applications (Zeng et al., 2022), verification (Huang et al., 2021; Thudi et al., 2022a), and security (Thudi et al., 2022b) of machine unlearning. However, research on these aspects is still in its early stages and is considered open problems (Xu et al., 2023b; Bourtoule et al., 2021). Although these aspects are undoubtedly important, our work is less related to these directions. As a result, we do not discuss these aspects in detail in this paper, but we acknowledge their significance and leave them for future research.
@@ 6. CONCLUSION
In this work, we study how to achieve complete, efficient unlearning without side effects on recommendation performance via the influence function. Towards this goal, we propose a novel and general unlearning method, which extends the influence function to estimate the spillover influence of unusable data to achieve more complete unlearning, and prunes less important parameters to further improve the unlearning efficiency. We conduct extensive experiments on two real-world datasets, providing insightful analysis for the effectiveness and great efficiency of our proposal in erasing interaction data. In the future, we will extend our method to deal with other types of data, e.g., sensitive features instead of interactions. Furthermore, we plan to apply our method to a diverse range of recommender models, such as factorization machines (Rendle, 2010; Zhang et al., 2023) and sequential models (Fang et al., 2020b), which can incorporate various side information. In particular, recommendation systems based on large language models have received significant attention recently (Wu et al., 2023b; Bao et al., 2023; Lin et al., 2023), and the massive parameters of such models pose new efficiency and efficacy challenges for existing unlearning approaches, including our proposed method. Hence, we will explore how to apply our method to these large models as well.

### v3

@@ Abstract.
In this work, we propose a new Influence Function-based Recommendation Unlearning (IFRU) framework, which efficiently updates the model without retraining by estimating the influence of the unusable data on the model via the influence function. In the light that recent recommender models use historical data for both the constructions of the optimization loss and the computational graph (e.g., neighborhood aggregation), IFRU jointly estimates the direct influence of unusable data on optimization loss and the spillover influence on the computational graph to pursue complete unlearning. Furthermore, we propose an importance-based pruning algorithm to reduce the cost of the influence function. IFRU is harmless and applicable to mainstream differentiable models. Extensive experiments demonstrate that IFRU achieves more than 250 times acceleration compared to retraining-based methods with recommendation performance comparable to full retraining. Codes are available at https://github.com/baiyimeng/IFRU.
Recommender System; Recommendation Unlearning; Privacy; Influence Function
copyright: acmlicensedjournal: TORSjournalyear: 2023journalvolume: xxxjournalnumber: xxxarticle: 1publicationmonth: 6doi: xxxxxx/xxxxxxccs: Information systems Recommender systemsccs: Information systems Collaborative filteringccs: Security and privacy Privacy protections
@@ 1. Introduction
We propose a new Influence Function-based Recommendation Unlearning (IFRU) framework to achieve more complete unlearning in recommendation. IFRU extends the influence function to quantify the influence of unusable data in the computational graph aspect, termed spillover influence. Specifically, IFRU expresses the computational graph change after erasing the unusable data in the optimization loss form, making spillover influence measurable like the direct influence. Taking one step further, we reduce the cost of influence function that computes the Hessian matrix with the size of model parameters by an importance-based pruning method to ignore less affected model parameters. IFRU avoids alterations to the training process and model architecture, emphasizing its capacity to maintain harmlessness while facilitating its application into mainstream models. We instantiate IFRU on two representative recommender models: MF (Koren et al., 2009) and LightGCN (He et al., 2020), and conduct experiments on two real-world datasets to demonstrate our method.
@@ 2.1. Problem Formulation
where $\hat{\theta}$ denotes the model parameters learned from $\mathcal{D}$, $L(\theta;\mathcal{D})$ denotes the total loss on $\mathcal{D}$, and $\ell(\cdot)$ is a recommendation loss function. In this study, our primary focus lies on the point-wise loss, particularly in investigating the Binary Cross-Entropy (BCE) loss. Nonetheless, our method can readily extend to other types of losses, such as pair-wise loss BPR (Rendle et al., 2009).
@@ 3. Methodology
In this section, we first introduce how IFRU achieves complete-efficient-harmless recommendation unlearning via the influence function (Section 3.1), and further acceleration via pruning (Section 3.2). We then present the instantiation of IFRU (Section 3.3) and further discussions of the proposal (Section 3.4).
@@ 3.1.1. Derivation of $\Delta$
Here, $\hat{\theta}_{\epsilon}$ represents the optimal model parameters under this scenario.
@@ 3.2. Unlearning with Pruning
Input: Pruning ratios $\{a_{0},...,a_{K}\}$, hyper-parameters $K$, training data $\mathcal{D}$, and the unusable data $\mathcal{D}_{r}$
1. For each $k\in\{0,1,\dots,K\}$, initialize $s_{v}^{k}$ as zero for each $v$ (a user or an item), and initialize $\mathcal{S}^{k}=\varnothing$ ;
@@ 3.3. Instantiation
Instantiation on LightGCN. For LightGCN, the computational graph for some remaining data will possibly change if removing $\mathcal{D}_{r}$, e.g., the aggregation operations could be different under $\mathcal{D}$ and $\mathcal{D}_{-r}$ as shown in Figure 5. That means, for some training data $(u,i,y_{ui})\in\mathcal{D}_{-r}$, we have $f_{\theta}(u,i;\mathcal{D})\neq f_{\theta}(u,i;\mathcal{D}_{-r})$. Thus $\mathcal{D}_{c}$ is not $\varnothing$. For simplicity, we set the number of graph convolution layers as $1$. In this case, it is easily verified that only the interaction (in $\mathcal{D}_{-r}$) of the users/items having interactions in $\mathcal{D}_{r}$ belongs to $\mathcal{D}_{c}$. That means:
@@ 3.4. Discussion
    As we consider both the spillover and direct influence, the estimated $\mathcal{I}(\hat{\theta};\mathcal{D}_{r})$ can faithfully reflect the true $\frac{d\hat{\theta}_{\epsilon}}{d\epsilon}\big{|}_{\epsilon=-\frac{1}{\mathcal{D}}}$. Then, if $|\mathcal{D}|$ is large enough, $o(\frac{1}{|\mathcal{D}|})$ is ignorable and $\hat{\theta}^{\prime}$ is a good approximation for the $\hat{\theta}^{*}$. Then, our unlearned model $f_{\hat{\theta}^{\prime}}(u,i;\mathcal{D}_{-r})$ can approach $f_{\hat{\theta}^{*}}(u,i;\mathcal{D}_{-r})$ well [^7] It is important to note that there are two underlying assumptions: 1) we assume that the errors caused by randomness and convergence in the learning models and calculating the influence can be ignored, and 2) it is assumed that the objective is strongly convex. Recent research by (Bae et al., 2022) suggests that the errors caused by randomness and convergence are very small. Furthermore, under non-convex cases, we can add the identity matrix with a damping term (Koh and Liang, 2017) to refine the defined influence, making it still applicable with smaller approximation errors (Bae et al., 2022; Warnecke et al., 2023).. Therefore, IFRU could achieve approximatively complete unlearning (defined in Section 2.1) if the size of training data is large enough. Regarding the pruning, empirical results (cf. Section 4.3.3) show that it will only slightly sacrifice the unlearning completeness since we only prune the model parameters less affected by erasing $\mathcal{D}_{r}$.
@@ 4.1.4. Hyper-parameters Setting
For fair comparisons, we use BCE loss (He et al., 2017) to learn model parameters for all methods. Meanwhile, we optimize all models using the Adam optimizer with a default batch size of 2048 and set the maximal number of epochs to 5000. The best hyper-parameters are determined based on the result of the validation set. The embeddings of the backbone recommender are initialized with the Gaussian distribution, the mean of which is fixed to zero and the standard deviation of which is tuned in the range of {1e-2, 1e-3, 1e-4}. The embedding size is tuned in the range of {32, 48, 64}. The early stopping strategy is applied to prevent the over-fitting problem, that is, terminating model training if the $AUC_{0}$ metric on the validation set does not increase in 50 successive epochs. The ratio of unusable data $\mathcal{D}_{r}$ to training data $\mathcal{D}$, i.e., $|\mathcal{D}_{r}|/|\mathcal{D}|$ is set to $2\%$ for MF and $1\%$ for LightGCN by default [^10] We would include a study that does not use HVP, where unlearning for LightGCN would be costly. Thus we take a smaller ratio for LightGCN to ensure the study could be easily conducted.. For all the compared methods, we tune the learning rate in the range of {1e-2, 1e-3, 1e-4}. For RecEraser, we adopt the Interaction-based Balanced Partition algorithm for data splitting and set the number of shards to 10, as suggested by the authors in their paper. The number of shards in SISA is also set to 10. For IFRU, we fix $K$ in Algorithm 1 to 1, as $1$-hop neighbors cover almost all users and items under the default ratio of unusable data. Then, we set the pruning ratios $(a_{0},a_{1})$ of Algorithm 1 to (1.0, 1.0) for MF and (1.0, 0.6) for LightGCN by default. The learning rate of Adam optimizer to calculate the defined influence is tuned in the range of {1e3, 2e3, 5e3, 1e4, 2e4, 5e4}. Notably, we have also tried smaller values for this learning rate and larger values for other learning rates. However, both of them lead to much worse results.
@@ 4.2.1. Accuracy Comparison
Harmless unlearning. Both incomplete unlearning and not achieving harmless unlearning can result in performance gaps between the Retrain method and unlearning methods, which may raise doubts about the superiority of IFRU in achieving harmless unlearning. To provide direct evidence for the superiority, we conduct an additional experiment. Specifically, after the backbone model was trained with a dataset, we forcibly set the unusable data $\mathcal{D}_{r}$ with an empty set and performed unlearning on it. In this case, if there is still a performance gap to Retrain, it must be due to not achieving harmless unlearning, as removing an empty set would not bring parameter changes for any method, including IFRU (seeing Equation (13)). The results of the experiment are summarized [^12] Notably, the training datasets for MF and LightGCN are $\mathcal{D}$ and $\mathcal{D}_{-r}$, respectively. Utilizing $\mathcal{D}_{-r}$ leads to performance discrepancies between Figure 4 and Table 2 for LightGCN. in Figure 4. The figure shows that IFRU/SCIF achieves the same performance as Original and Retrain, confirming that unlearning with influence function can help achieve harmless unlearning. In contrast, RecEraser and SISA show performance gaps to Original and Retrain, indicating that harmless unlearning is not achieved.
@@ 5.1. Machine Unlearning
In addition to researching how to achieve unlearning, there are also some works that specifically focus on the benchmarks (Schelter, 2020), applications (Zeng et al., 2022), verification (Huang et al., 2021; Thudi et al., 2022a), and security (Thudi et al., 2022b) of machine unlearning. However, research on these aspects is still in its early stages and is considered open problems (Xu et al., 2023b; Bourtoule et al., 2021). Although these aspects are undoubtedly important, our work is less related to these directions. As a result, we do not discuss these aspects in detail in this paper, but we acknowledge their significance and leave them for future research.
@@ 6. Conclusion
In this work, we study how to achieve complete, efficient unlearning without side effects on recommendation performance via the influence function. Towards this goal, we propose a novel and general unlearning method, which extends the influence function to estimate the spillover influence of unusable data to achieve more complete unlearning, and prunes less important parameters to further improve the unlearning efficiency. We conduct extensive experiments on two real-world datasets, providing insightful analysis for the effectiveness and great efficiency of our proposal in erasing interaction data. In the future, we will extend our method to deal with other types of data, e.g., sensitive features instead of interactions. Furthermore, we plan to apply our method to a diverse range of recommender models, such as factorization machines (Rendle, 2010; Zhang et al., 2023) and sequential models (Fang et al., 2020b), which can incorporate various side information. In particular, recommendation systems based on large language models have received significant attention recently (Wu et al., 2023b; Bao et al., 2023; Lin et al., 2023), and the massive parameters of such models pose new efficiency and efficacy challenges for existing unlearning approaches (Hu et al., 2024), including our proposed method. Hence, we will explore how to apply our method to these large models as well.